# Learning to Act without Actions

### v1

@@ Abstract
Pre-training large models on vast amounts of web data has proven to be an effective approach for obtaining powerful, general models in several domains, including language and vision. However, this paradigm has not yet taken hold in deep reinforcement learning (RL). This gap is due to the fact that the most abundant form of embodied behavioral data on the web consists of videos, which do not include the action labels required by existing methods for training policies from offline data. We introduce Latent Action Policies from Observation (LAPO), a method to infer latent actions and, consequently, latent-action policies purely from action-free demonstrations. Our experiments on challenging procedurally-generated environments show that LAPO can act as an effective pre-training method to obtain RL policies that can then be rapidly fine-tuned to expert-level performance. Our approach serves as a key stepping stone to enabling the pre-training of powerful, generalist RL models on the vast amounts of action-free demonstrations readily available on the web.
@@ 1 Introduction
Training on web-scale data has shown to be an effective approach for obtaining powerful models with broad capabilities in domains including language and vision (Radford et al., 2019; Caron et al., 2021). Much recent work has thus sought to apply the same paradigm in deep reinforcement learning (RL, Sutton & Barto, 2018), in order to learn generalist policies from large amounts of web data (Baker et al., 2022; Reed et al., 2022). However, common methods for learning policies from offline demonstrations, such as imitation learning (Pomerleau, 1988; Ross & Bagnell, 2010) and offline RL (Kumar et al., 2020; Levine et al., 2020), generally require data to include action or reward labels, which are missing from observational data, such as videos found on the web.
@@
In this work we introduce Latent Action Policies from Observation (LAPO), a method for learning useful, pre-trained policies from purely observational data. LAPO is founded on the key insights that (1) some notion of a latent action that explains each environment transition can still be inferred from observations alone, and (2) given such inferred latent actions per transition, a latent-action policy can be obtained using standard imitation learning methods. Our experiments provide strong evidence that such latent policies accurately capture the observed behavior of the expert by showing that the latent policy can be rapidly finetuned to recover expert-level performance in the observed tasks. This fact makes the approach underlying LAPO a significant stepping stone toward pre-training general, rapidly-adaptable policies on massive action-free demonstration datasets, such as the vast quantities of videos available on the web.
@@
Crucially, LAPO learns to infer latent actions, and consequently, obtain latent action policies in a fully unsupervised manner. LAPO is similar to prior work (Torabi et al., 2018; Schmeckpeper et al., 2020; Baker et al., 2022; Zheng et al., 2023) in that we first train an inverse dynamics model (IDM) that predicts the action taken between two consecutive observations, and then use this IDM to add action labels to a large action-free dataset. However, unlike these prior works, which rely on some amount of ground-truth action-labelled data for training the IDM, LAPO does not make use of any labels and infers latent action information purely from observed environment dynamics. To do so, the IDM in LAPO learns to predict latent actions rather than true actions. These latent actions take the form of a learned representation that explains an observed transition. In order to learn this representation, LAPO makes use of a simple unsupervised objective that seeks to establish predictive consistency between the IDM and a separate forward dynamics model (FDM). The FDM is trained on transitions consisting of two consecutive observations $(o_{t},o_{t+1})$ to predict the future observation $o_{t+1}$, given the past observation $o_{t}$ and a latent action. The latter is generated by the IDM after observing both past and future observations. Thus, unlike the FDM which sees only the past, the IDM has access to both the past and future and learns to pass useful information about the future to the FDM through the latent action. By making the latent action an information bottleneck (Tishby et al., 2000), we prevent the IDM from simply forwarding the entire future observation to the FDM and instead force it to learn a highly compressed encoding of state transitions. We refer to this encoding as a latent action, as it captures the visible effects of the agent’s true actions.
@@
Figure 1: UMAP projection of the learned latent action space for Miner alongside sample next-state predictions generated by the FDM for each cluster of latent actions. Each point represents a continuous (i.e. pre-quantization) latent action generated by the IDM for a transition in the observation-only dataset. Each point is color-coded by the true action taken by the agent at that transition. For clarity, NOOP actions are omitted. The structure of the latent action space is highly interpretable and closely corresponds to the true action space, even though no ground-truth action labels were used during training.
@@
Our experiments in Section 6 evaluate the potential of LAPO as a method for pre-training useful RL policies from action-free demonstrations. First, we train a latent IDM via LAPO on large expert-level action-free offline datasets for each of the 16 games of the Procgen Benchmark (Cobbe et al., 2019; 2020). We observe that the structure of the learned latent action spaces are highly interpretable, often exhibiting clusters that closely align with the true action space (see Figure 1). This is a remarkable result, as our approach uses zero ground-truth action labels and instead recovers all action information purely from observed dynamics. Next, we leverage the learned latent representation for downstream online RL tasks. For this, we first use the IDM to assign latent action labels to each transition in the same observation-only dataset on which it was trained. We then perform behavior cloning on the resulting action-observation dataset to learn a latent-action policy, which seeks to imitate the latent actions predicted by the IDM. Finally, we seek to transfer the latent action policy to the true action space, in order to deploy and evaluate it in the online environment. We show that simply fine-tuning the last layers of the latent policy in the original environment with a general-purpose RL algorithm allows the policy to rapidly adapt to the true action space and recover (or at times exceed) expert performance. Further, we show that we can alternatively train a latent to true action decoder on a small labeled dataset (on the order of a few hundred transitions). This action decoder, when composed with the latent action policy, attains significantly better performance than training PPO from scratch with no need for environment interaction or reward labels. This indicates that latent policies produced by LAPO capture meaningful behavior in the latent action space.
In summary, we show that a simple unsupervised objective can be used to recover latent action information and learn latent policies purely from observed environment dynamics. Our experiments across challenging procedurally-generated environments show that the latent representations learned by our method capture much of the structure underlying the true action space. We show that the resulting latent-space policies can be rapidly adapted to recover expert-level performance in the true action space. We believe our approach can serve as a key stepping stone towards training powerful, generalist policies on web-scale video-based demonstrations.
@@ 2 Related work
Most similar to our approach, ILPO (Edwards et al., 2019) also aims to infer latent actions via a dynamics modelling objective. However, unlike LAPO, ILPO learns a latent policy rather than an IDM jointly with the FDM, and uses a discrete rather than a continuous latent action space. For each FDM update, ILPO identifies the discrete latent action that minimizes the next-state prediction error, and only optimizes the FDM for that specific action. In practice, this objective can lead to mode collapse, where a feedback loop causes only a small number of (or only one) discrete latent to ever be selected (Struckmeier & Kyrki, 2023). By using an IDM rather than a discrete latent policy, LAPO avoids enumerating all hypothetically possible transitions and instead learns the latent action corresponding to the actual observed transition. To train the policy, ILPO minimizes the difference between the true next state and the expectation of the next state under the policy. This loss is ill-conditioned, as even when it is minimized, the next-state predictions for individual latent actions do not have to align with the environment. Moreover, LAPO’s use of continuous rather than finite discrete latents may allow better modeling of more complex environment dynamics, as any information useful for prediction can be captured in the latent action representation. Further, ILPO’s use of discrete latent actions makes its computational complexity linear rather than constant in the size of the latent action space, due to the required enumeration of all actions.
@@
LAPO differs from previous methods for imitation learning from observation (IfO, Torabi et al., 2019a; b; Yang et al., 2019), which typically require the imitating policy to have access to the true action space when training on action-free demonstrations. Crucially, unlike these prior approaches, LAPO does not require access to the ground-truth action space to learn from action-free demonstrations. Other methods leverage observation-only demonstrations for purposes other than learning useful behaviors. Intention-Conditioned Value Functions (ICVF Ghosh et al., 2023) uses a temporal difference learning objective based on observation-only demonstrations to learn state representations useful for downstream RL tasks. Aytar et al. (2018) propose to use expert demonstrations during online RL to guide the agent along the expert’s trajectories. For this, they propose two auxiliary classification tasks for learning state representations based on observation-only demonstrations.
@@ 3.4 Vector-quantization
4 Latent Action Policies from Observation
@@ 4.3 Decoding Latent Actions
The policy $\pi$, obtained via BC, produces actions in a continuous latent space. We consider two ways to adapt $\pi$ to produce outputs in the true action space of the online RL environment.
Online RL policy fine-tuning
When given access to the online RL environment, we can simply finetune the latent action policy in the environment using RL. This approach, which we refer to as online RL-fine-tuning, treats the latent behavior cloning step as a pretraining phase. During RL fine-tuning, the policy only needs to be adapted slightly to output actions in the true action space, e.g. by replacing the latent action policy head with a true action policy head. Moveover, the performance of the policy can potentially be improved beyond that of the data-generating policy. In particular, when the data was generated by a mixture of policies of different skill levels or pursuing different goals (as is the case with data from the web), this step can extract the data-generating policies more closely aligned with the reward function used for fine-tuning. Our experiments make use of a single data-generating expert, which can exhibit a similar diversity of outcomes across trajectories.
Offline action decoding
If only offline, action-labeled transitions are available, we can train a small decoder network $d\colon\mathcal{Z}\rightarrow\mathcal{A}$, mapping the predicted latent action for each labeled transition to the ground-truth action. This decoder can be composed with the latent policy and the resulting decoded latent policy $d\circ\pi$ can be deployed in the online environment.
@@ 5 Experimental Setting
We use the IMPALA-CNN introduced by Espeholt et al. (2018) to implement both our policy and IDM with a 4x channel multiplier as used by Cobbe et al. (2020), and U-Net (Ronneberger et al., 2015) based on a ResNet backbone (He et al., 2016) with approximately 8M parameters for the FDM. In the offline action decoding setting, the decoder is a fully-connected network with hidden sizes $(128,128)$. We use an EMA-based update (Polyak & Juditsky, 1992) for the vector quantization embedding. We use a single observation of additional pre-transition context, i.e. $k=1$.
We use Proximal Policy Optimization (PPO, Schulman et al., 2017) for online fine-tuning and initialize the policy with the weights of the latent policy, except for the last layer, which is re-initialized with output dimensionality $|\mathcal{A}|$ instead of $|\mathcal{Z}|$. We keep all convolutional layers frozen and only fine-tune the actor and policy heads as well as the previous fully-connected layer. We found that a much larger learning rate of 0.01 can be stably used when only fine-tuning these three layers. We similarly tuned the learning rate for training the full network from scratch, but found no improvement compared to the original value of 5e-4. Other hyperparameters are given in Appendix A.5.
@@ 6 Results and Discussion
6.1 Online RL policy fine-tuning
Figure 3: Left: Mean test returns over the course of training for PPO fine-tuning of a pretrained LAPO policy and PPO from scratch (averaged across 3 seeds). Right: Mean test returns relative to per-environment expert policies averaged across all 16 Procgen environments. Error bars indicate standard deviation across seeds.
We now discuss our results for fine-tuning the latent policy in the online environment using RL. Recall, that this is the third stage of our approach, after first training a latent IDM and then obtaining a latent policy via behavior cloning. In this experiment, we investigate how LAPO’s pretraining step is able to speed up online RL compared to learning a policy from scratch. Since our work focuses on the setting where large amounts of observation-only data are freely available, but environment interaction is limited or expensive, we run PPO for 4M steps rather than the 25M steps suggested by Cobbe et al. (2020). Experiments are performed separately per environment. As can be seen in Figure 3, fine-tuning policies pretrained with LAPO recovers expert-level performance within only 4M frames, while PPO from scratch reaches only 44% of expert performance in the same period. We highlight that in 8 of 16 games, we are able to reach 80% of expert performance within only 300k frames, and in 3 of 16 games, our approach exceeds expert performance within 4M frames. We also provide results for an ablation of our method that does not apply vector-quantization to latent actions, showing that VQ is an important component of our method.
We use ILPO as our primary baseline, however we found that its policy immediately collapses in several Procgen tasks. In tasks where it does not collapse, online fine-tuning of the ILPO policy did not perform better than learning from scratch. While Edwards et al. (2019) do provide positive results for the task CoinRun (which we reproduce in Appendix A.4), their experiments target only a single level of the environment. We thus hypothesize that ILPO’s policy collapse is in part due to the modeling challenge posed by the more visually diverse levels in the fully procedurally-generated setting. By using an IDM rather than a policy plus a high-dimensional, continuous latent action space, LAPO can better model stochasticity (such as visual noise) and unobserved information (such as off-screen parts of the game). In contrast, it is likely difficult for ILPO to model the full distribution of possible transitions using only a small set of discrete latents.
6.2 Offline action decoding
@@
Next, we consider an offline approach to training a latent action decoder. Here, we simply train a decoder network that maps latent actions to true actions using a small amount of action-labeled demonstration data. We then compose the decoder with the latent policy and evaluate this decoded latent policy in the environment. As we will see in Section 6.3, the latent and true action spaces share similar structure. This similarity allows an effective decoder to be trained on only a miniscule amount of action-labeled data. Figure 4 shows that an offline decoder trained on less than 256 labeled transitions achieves the same performance as a policy trained from scratch for 4M steps with PPO. However, we observe that with increasing dataset size, performance eventually plateaus below the level of RL online fine-tuning. This is likely because, as previously noted, latent actions are not necessarily state-invariant. A decoder that only has access to a latent action, but not the state in which that action is being taken, is not always able to successfully decode it. We provide per-environment results for offline decoding in Appendix A.1.

### v2

@@ Abstract
Pre-training large models on vast amounts of web data has proven to be an effective approach for obtaining powerful, general models in domains such as language and vision. However, this paradigm has not yet taken hold in reinforcement learning. This is because videos, the most abundant form of embodied behavioral data on the web, lack the action labels required by existing methods for imitating behavior from demonstrations. We introduce Latent Action Policies (LAPO), a method for recovering latent action information—and thereby latent-action policies, world models, and inverse dynamics models—purely from videos. LAPO is the first method able to recover the structure of the true action space just from observed dynamics, even in challenging procedurally-generated environments. LAPO enables training latent-action policies that can be rapidly fine-tuned into expert-level policies, either offline using a small action-labeled dataset, or online with rewards. LAPO takes a first step towards pre-training powerful, generalist policies and world models on the vast amounts of videos readily available on the web. Our code is available here: https://github.com/schmidtdominik/LAPO.
@@ 1 Introduction
Training on web-scale data has shown to be an effective approach for obtaining powerful models with broad capabilities in domains including language and vision (Radford et al., 2019; Caron et al., 2021). Much recent work has thus sought to apply the same paradigm in deep reinforcement learning (RL, Sutton & Barto, 2018), in order to learn generalist policies from massive amounts of web data (Baker et al., 2022; Reed et al., 2022). However, common methods for learning policies from offline demonstrations, such as imitation learning (Pomerleau, 1988; Ross & Bagnell, 2010) and offline RL (Kumar et al., 2020; Levine et al., 2020), typically require action or reward labels, which are missing from purely observational data, such as videos found on the web.
In this work we introduce Latent Action Policies (LAPO), a method for recovering latent-action information from videos. In addition to inferring the structure of the underlying action space purely from observed dynamics, LAPO produces latent-action versions of useful RL models including forward-dynamics models (world models), inverse-dynamics models, and importantly, policies.
@@
LAPO is founded on the key insights that (1) some notion of a latent action that explains an environment transition can be inferred from observations alone, and (2) given such inferred latent actions per transition, a latent-action policy can be obtained using standard imitation learning methods. Our experiments provide strong evidence that such latent policies accurately capture the observed expert’s behavior, by showing that they can be efficiently fine-tuned into expert-level policies in the true action space. This results makes LAPO a significant step toward pre-training general, rapidly-adaptable policies on massive action-free demonstration datasets, such as the vast quantities of videos available on the web.
@@
Crucially, LAPO learns to infer latent actions, and consequently, obtain latent action policies in a fully unsupervised manner. LAPO is similar to prior work (Torabi et al., 2018; Schmeckpeper et al., 2020; Baker et al., 2022; Zheng et al., 2023) in that we first train an inverse dynamics model (IDM) that predicts the action taken between two consecutive observations, and then use this IDM to add action labels to a large action-free dataset. However, unlike these prior works, which rely on some amount of ground-truth action-labelled data for training the IDM, LAPO does not make use of any labels and infers latent action information purely from observed environment dynamics. To do so, the IDM in LAPO learns to predict latent actions rather than true actions. These latent actions take the form of a learned representation that explains an observed transition. To learn this representation, LAPO makes use of a simple unsupervised objective that seeks to establish predictive consistency between the IDM and a separate forward dynamics model (FDM). The FDM is trained on transitions consisting of two consecutive observations $(o_{t},o_{t+1})$ to predict the future observation $o_{t+1}$, given the past observation $o_{t}$ and a latent action. This latent action is generated by the IDM given both past and future observations. Thus, unlike the FDM which sees only the past, the IDM has access to both the past and future and learns to pass useful information about the future to the FDM through the latent action. By making the latent action an information bottleneck (Tishby et al., 2000), we prevent the IDM from simply forwarding the entire future observation to the FDM, thereby forcing it to learn a highly compressed encoding of state transitions. We refer to this encoding as a latent action, as it captures the observable effects of the agent’s true actions.
@@
Figure 1: UMAP projection of the learned latent action space for Miner alongside illustrative next-state predictions generated by the FDM for each cluster of latent actions. Each point represents a continuous latent action generated by the IDM for a transition in the video dataset. Each point is color-coded by the true action taken by the agent at that transition. For clarity, NOOP actions are omitted. The structure of the latent action space is highly interpretable and closely corresponds to the true action space, even though no ground-truth action labels were used during training.
@@
In our experiments in Section 6, we train a latent IDM via LAPO on large expert-level action-free offline datasets for each of the 16 games of the Procgen Benchmark (Cobbe et al., 2019; 2020). We observe that the structure of the learned latent action spaces are highly interpretable, often exhibiting clusters that closely align with the true action space (see Figure 1). This is a remarkable result, as our approach uses zero ground-truth action labels and instead recovers all action information purely from observed dynamics. To go beyond merely recovering action information, we then show how we can leverage the learned latent representation to quickly obtain powerful expert-level policies in downstream tasks: We first use the IDM to assign latent action labels to each transition in the same video dataset on which it was trained. We then perform behavior cloning on the resulting action-observation dataset to learn a latent-action policy, which imitates the latent actions predicted by the IDM. Finally, to deploy the policy in the online environment, we seek to turn it from a latent action policy into a policy in the true action space. We show that if a tiny, action-labeled dataset is available, a latent $\rightarrow$ true action decoder can be learned in a supervised manner and used to decode the outputs of the latent policy. This approach is extremely data-efficient, requiring only $\sim$200 labeled transitions to exceed the performance of 4M steps of regular RL via PPO. Alternatively, if an online environment is available, we show that simply fine-tuning the last layers of the latent policy with a general-purpose RL algorithm allows the policy to rapidly adapt to the true action space and recover (or at times exceed) expert performance. These results indicate that latent policies produced by LAPO capture meaningful behavior in the latent action space.
While we take a first step in demonstrating the potential of these latent representations for obtaining powerful policies in Section 6, the fundamental contribution of this work is simply showing that comprehensive action information can be recovered from pure video through unsupervised objectives, such as the one proposed herein. This result opens the door to future work that employs representation learning methods like LAPO to train powerful, generalist policies and world models on massive video datasets from the web.
@@ 2 Related work
Most closely related to our approach, ILPO (Edwards et al., 2019) aims to infer latent actions via a dynamics modelling objective. However, unlike LAPO, ILPO learns a latent policy rather than an IDM jointly with the FDM, and uses a discrete rather than a continuous latent action space. For each FDM update, ILPO identifies the discrete latent action that minimizes the next-state prediction error, and only optimizes the FDM for that specific action. In practice, this objective can lead to mode collapse, where a feedback loop causes only a small number of discrete latents to ever be selected (Struckmeier & Kyrki, 2023). By using an IDM rather than a discrete latent policy, LAPO avoids enumerating all hypothetically possible transitions and instead learns the latent action corresponding to the actual observed transition. To train the policy, ILPO minimizes the difference between the true next state and the expectation of the next state under the policy. This loss is ill-conditioned, as even when it is minimized, the next-state predictions for individual latent actions do not have to align with the environment. Moreover, LAPO’s use of continuous rather than finite discrete latents may allow better modeling in complex and partially-observable environments, as any information useful for prediction can be captured within the latent representation. We provide evidence for this in Appendix A.5, where we show that ILPO breaks down when applied to more than a single Procgen level at a time. Finally, ILPO’s use of discrete latents makes its computational complexity linear rather than constant in the latent dimensionality, due to the required enumeration of all actions.
@@
LAPO differs from previous methods for imitation learning from observation (IfO; Torabi et al., 2019a; b; Yang et al., 2019), which typically require the imitating policy to have access to the true action space when training on action-free demonstrations. Crucially, unlike these prior approaches, LAPO does not require access to the ground-truth action space to learn from action-free demonstrations. Other methods leverage observation-only demonstrations for purposes other than learning useful behaviors. Intention-Conditioned Value Functions (ICVF; Ghosh et al., 2023) uses a temporal difference learning objective based on observation-only demonstrations to learn state representations useful for downstream RL tasks. Aytar et al. (2018) propose to use expert demonstrations during online RL to guide the agent along the expert’s trajectories. For this, they propose two auxiliary classification tasks for learning state representations based on observation-only demonstrations.
@@ 3.4 Vector-quantization
4 Latent Action Policies
@@ 4.3 Decoding Latent Actions
The policy $\pi$, obtained via BC, produces actions in the latent action space. We now consider different ways to adapt $\pi$ to produce outputs in the true action space, depending on the kinds of data available.
Small action-labeled dataset.
If a small dataset of true-action-labeled transitions is available, we can train an action decoder, $d\colon\mathcal{Z}\rightarrow\mathcal{A}$, to map the IDM-predicted latent action for each transition to the known ground-truth action. The trained decoder can then be composed with the frozen latent policy and the resulting decoded latent policy, $d\circ\pi\colon\mathcal{O}\rightarrow\mathcal{A}$, can be deployed online.
Online environment.
By interacting with the online environment we get both action-labeled transitions, which can be used to train a decoder in a supervised manner as described above, and a reward signal, which can be used for online RL. By doing both simultaneously, we can quickly bootstrap a crude, initial decoder via supervised learning, and then fine-tune the entire composed policy $d\circ\pi$ by directly optimizing for the environment’s reward function via RL. Through RL, the performance of the policy can be potentially be improved beyond that of the data-generating policy. In particular, when the data was generated by a mixture of policies of different skill levels or pursuing different goals (as is the case with data from the web), this step can extract the data-generating policies more closely aligned with the reward function used for fine-tuning.
@@ 5 Experimental Setting
We use the IMPALA-CNN (Espeholt et al., 2018) to implement both our policy and IDM with a 4x channel multiplier as used by Cobbe et al. (2020), and U-Net (Ronneberger et al., 2015) based on a ResNet backbone (He et al., 2016) with approximately 8M parameters for the FDM. The latent action decoder is a fully-connected network with hidden sizes $(128,128)$. We use an EMA-based update (Polyak & Juditsky, 1992) for the vector quantization embedding. We use a single observation of additional pre-transition context, i.e. $k=1$. When decoding the latent policy in the online environment, we have two decoder heads on top of the latent policy whose logits are averaged per action before applying softmax. One head is trained in a supervised manner on $(a_{t},z_{t})$ tuples from historical data, the other is trained via Proximal Policy Optimization (PPO, Schulman et al., 2017). We keep all convolutional layers frozen and found that a much larger learning rate of 0.01 can be stably used when only training these final few layers. We similarly tuned the learning rate for training the full network from scratch, but found no improvement compared to the original value of 5e-4. Other hyperparameters are given in Appendix A.4.
@@ 6 Results and Discussion
6.1 Decoding the latent policy online
Figure 3: Left: Mean episodic returns (over the course of training) for decoding LAPO’s latent policy and PPO from scratch (averaged across 3 seeds). Right: Mean test returns relative to per-environment expert policies averaged across all 16 Procgen environments. Error bars indicate standard deviation across seeds.
We now discuss our results for decoding the latent policy in the online environment. Recall, that this is the third stage of our approach, after first training a latent IDM, followed by obtaining a latent policy via behavior cloning. In this experiment, we compare how decoding the latent policy (via supervised learning and RL) compares to learning a policy from scratch purely from environment interaction via PPO. Since our work focuses on the setting where large amounts of observation-only data are freely available, but environment interaction is limited or expensive, we run PPO for 4M steps rather than the 25M steps suggested by Cobbe et al. (2020). Experiments are performed separately per environment. As can be seen in Figure 3, using LAPO’s latent policy we are able to fully recover expert performance within only 4M frames, while PPO from scratch reaches only 44% of expert performance in the same period. We highlight that in 9 of 16 tasks, our approach exceeds the performance of the expert (which was trained for 25M frames) within only 4M frames. We also provide results for two ablations. The “RL, no-SL” ablation trains the decoder only using reinforcement learning. In this setting, LAPO can be viewed as an RL-pretraining step that produces an initial policy representation that is useful for downstream RL. The “RL, no-SL, no-VQ” ablation additionally removes vector-quantization, showing that VQ is an important component of our method.
We use ILPO as our primary baseline, but found that its policy immediately collapses in several Procgen tasks. When it does not collapse, online decoding of the ILPO policy did not perform better than PPO from scratch. While Edwards et al. (2019) do provide positive results for CoinRun (which we reproduce in Appendix A.5), their experiments target only individual levels of this single task. We thus hypothesize that ILPO’s policy collapse is in part due to the modeling challenge posed by the visually diverse levels in the full, procedurally-generated setting. By using an IDM rather than a policy, plus a high-dimensional, continuous latent space, LAPO can capture stochasticity and unobserved information (such as off-screen parts of the game). In contrast, it is likely difficult for ILPO to model the full distribution of possible transitions using only a few discrete latents.
6.2 Decoding the latent policy offline
@@
Next, we consider training a latent action decoder fully-offline using only a small action-labeled dataset and no access to the online environment. The trained decoder is then composed with the latent policy to map the policy’s actions into the true action space. As we will see in Section 6.3, the latent and true action spaces share similar structure. This similarity allows an effective decoder to be trained on only a miniscule amount of action-labeled data. Figure 4 shows that a decoder trained on less than 256 labeled transitions matches the performance of a policy trained from scratch for 4M steps. However, we observe that with increasing dataset size, performance eventually plateaus below the level of online decoder training. This is likely because, as previously noted, latent actions are not necessarily state-invariant. A decoder that only has access to a latent action, but not the state in which that action is being taken, may not always be able to successfully decode it. We provide per-environment results in Appendix A.1.